OpenStudy (anonymous):

A roll of paper towels is wound around a hollow cardboard tube. The cardboard tube in the middle of the roll has an outside radius of 2.36 centimeters. The thickness of the paper is 0.04 centimeters. The sequence of distances of the loops of paper away from the center of the roll, in centimeters, is the following: a1, a2, a3, a4, a5, ... = 2.40, 2.44, 2.48, 2.52, 2.56, ... What is the radius of the 95th loop of paper, starting from the center of the tube? A. 6.16 cm b. 6.36 cm c. 6.26 cm d. 6.46 cm

OpenStudy (anonymous):

This is an arithmetic sequence with a common difference of .04. Any term, An = A1 + (n - 1)d where A1 is the first term, n is the index of (in our case 95) and d is the common difference. A_95 = 2.40 + (95 - 1) * 0.04 = 6.16

OpenStudy (anonymous):

n is the index of the term*